### v1

@@ 2 Model Definitions
    The trajectory terminates upon reaching $A_{\omega}$ or when no such single tile configuration $a_{t}$ can be added. In the latter case, the trajectory is said to have been prematurely terminated.
@@ 3 Results
3.1 The Backboned aTAM admits a Universal Assembly Kit
Our results are concerned with finding, for both the backboned and sequenced aTAM, a finite tile type set and corresponding strength function that allows the deterministic assembly of any shape within the shape space of the model. Here, the starting configuration must be empty set and the sequence of tiles can be arbitrary, within those sequences permitted by the finite tile set. We call such a tile type set a universal assembly kit. We begin our results with the following Lemma.
@@ 3.1 The Backboned aTAM admits a Universal Assembly Kit
Proof 3.1.
This lemma trivially proceeds from definition 2, as added tiles must neighbour the last added tile.
@@
Theorem 3.2.
A universal assembly kit for the backboned aTAM exists.
Proof 3.3.
Since any shape that can be assembled by a backboned aTAM system is a shape of some self-avoiding walk, we can always assemble any shape (achievable by a backboned aTAM system) by encoding the set of left, right or forward moves of the underlying self-avoiding walk. We use glues $0,1$ and $2$, along with the following strength function:
[…]
Then, each left, right or forward move can be performed by a specific “directed” tile, shown in Figure 3. An additional tile encoding the start of the shape is required, but there is no need for an end tile as any of the directed tiles can be placed at the end of any self-avoiding walk without impacting the final shape. Hence, a universal assembly kit with 4 tiles is sufficient to assemble any shape achievable by the backboned aTAM, without requiring a starting seed configuration or needing to scale up the shape.
Given this theorem, the following corollary trivially holds.
Corollary 3.4.
Proof 3.5.
Any self-avoiding path can be constructed using the method in Theorem 3.2.
These initial results are fairly straightforward. However, the construct we have arrived at is somewhat artificial when considering the physical system being emulated (co-translational folding), since the shape is completely determined by the backbone routing and non-backbone-adjacent interactions are irrelevant. This result is only possible due to the use of neutral interactions, or by allowing adjacent tiles to have weak repulsive interactions (that are not sufficiently strong to prevent the addition of a tile). We now consider, therefore, whether the existence of a universal assembly kit is maintained even if we require that all inter-tile interactions in any configuration are attractive, so that neighbours that are brought together by the folded pathway must interact attractively.
Theorem 3.6.
@@
Proof 3.7.
The existence of some infinite tile set is trivial, as in the construction of all possible shapes, one can use the tile set in Theorem 3.2, but replace all glue types of all faces that neighbour another tile face with some unique attractive glue type. However,a universal assembly kit requires a finite number of tiles and reusing the same attractive glue type multiples times within a configuration can result in non-deterministic assembly as tiles can be attracted into incorrect positions (Figure 4). Hence, we must proceed to construct a scheme for numbering tile types such that only a finite number of glue types $N$ are required while still guaranteeing deterministic shape assembly. Let the glue types be labelled $\{0,3,4,5...N-1\}$, and let the strength function $g$ be defined as follows.
[…]
@@
These tile types are constructed by analogy with the tile types used in the proof of Theorem 3.2. For example, $\hat{\Theta}_{0}$ is a set of tile types with glue types 1 and 2 (backbone glue types) in the same pattern as tile type 0 in Fig. 3, but with variable glues on the other faces; $\hat{\Theta}_{0}(\sigma,\sigma^{\prime})$ identifies a specific member of that set. Non-backbone glue types are divided into zero and non-zero types; we denote by $B^{c}({\Theta})$ the set of non-zero, non-backbone glue types on tile type ${\Theta}$. Tile faces endowed with backbone and non-backbone glue types are correspondingly called backbone faces and non-backbone faces, respectively.
Figure 4: A figure illustrating the difficulties associated with assembly without neutral interactions. While in principle one can replace the neutral interface type 00 with a self-attractive interface type $3$, such attractive interfaces can pull tiles towards unintended positions (right).
@@
We present an algorithm for selecting a sequence $Q$, using a subset of the tiles from Definition 3.8 with $N=7$, to define a backboned aTAM system $(\mathcal{A},Q,g)$ that can deterministically assemble any given shape within the shape space of the backboned aTAM. The interacting directed tile types in Definition 3.8, along with the strength function in Equation 2, therefore define a universal assembly kit for the backboned aTAM, proving the existence of such a kit by construction.
Let $\Psi=(A_{\rm empty},A_{1},...A_{\omega})$ be a complete trajectory that assembles a shape $\overline{S}$, defining an arbitrary Hamiltonian path through the shape. A sequence of backbone faces consistent with this trajectory can then be selected as in the proof of Theorem 3.2, fixing whether each tile in the sequence is drawn from the set $\hat{\Theta}_{T}$, $\hat{\Theta}_{0}$, $\hat{\Theta}_{1}$, $\hat{\Theta}_{3}$ or $\hat{\Theta}_{H}$. Hence, we only need to select the non-backbone interactions of tile types in the sequence in such a way as to ensure deterministic production of the desired shape.
@@
    Non-backbone faces of $a_{t}$ with no neighbour in $A_{\omega}$ (the terminal configuration of $\Psi$) are assigned the repulsive glue type 0.
@@
    Non-backbone faces of $a_{t}$ with no neighbour in $A_{t}$ but with a neighbour in $A_{\omega}$ are assigned an unknown glue type $\sigma_{u}$
@@
Figure 6: An illustration of the inequalities on the faces of tiles when constructing an assembly kit for the backboned aTAM without neutral interactions. After setting the glue types of known faces, there remains one face with unknown glue type $\sigma_{u}$. $\sigma_{u}\neq 3$, $\sigma_{u}\neq 5$ and $\sigma_{u}\neq 6$, so $\sigma_{u}=4$ is the only correct option if we restrict ourselves to $4$ attractive non-backbone glue types.
@@
The arguments above break down for the penultimate tile as the final tile can have three non-backbone-connected neighbours. However, in this case all of the final neighbours of $a_{\omega}$ and $a_{\omega-1}$ are already in place. Therefore the non-backbone glue types of $a_{\omega-1}$ are either specified by neighbouring tiles that are present in the configuration $A_{\omega-1}$, or 0. They cannot, therefore, cause any ambiguity when $a_{\omega}$ is placed, with glue types set either to match neighbours present in configuration $A_{\omega}$, or 0. Taking $N=7$, 208 tiles are defined by Definition 3.8. This number provides an upper bound on the size of the minimal universal assembly kit for the backboned aTAM.
This result again contrasts with the base aTAM, for which a universal assembly kit does not exist. Hence, in some sense, backbones increase our ability to reliably assemble a complex structure from a small set of subunits.
3.2 Tile Complexity Comparisons between the Backboned and Base aTAM
As well as considering universal assembly kits, we may also ask whether the backboned aTAM requires fewer unique tiles (that is, has a lower tile complexity) to deterministically produce a single target shape than the base temperature-1 aTAM. For large enough target shapes, the argument in Section 3.1 suggests that the backboned aTAM will incur a smaller tile complexity, as the tile complexity for a specific target is upper-bounded by the universal kit of size 208, whereas we expect tile complexity to grow indefinitely with the target shape size for the base aTAM ([24]).
@@ 3.2 Tile Complexity Comparisons between the Backboned and Base aTAM
Figure 7: Backbones can reduce the minimal assembly kit for single target assembly, even at relatively low system size. We consider the construction of rectangles (a) and bulged rectangles (b) with $m$ rows and $n$ columns. The backbone path for the backboned aTAM is given in purple. The bulge location is fixed to the second and third rows of the rightmost column. The number of tile types required in the backboned aTAM for either shape peaks at $14\ll 208$, the upper bound on minimal universal assembly kit size. We plot the difference in upper bound estimates of assembly kit sizes required to construct rectangles (c) and bulged rectangles (d) using the base aTAM (tile complexity denoted by $K_{u}$) and the backboned aTAM (tile complexity denoted by $K_{b}$) assembly systems.
@@
We can apply the method in ([1]) to upper bound a minimal tile complexity for a given shape for the base aTAM, and we can apply the algorithm constructed in Theorem 3.6 to find an upper bound on the tile complexity required to assemble the same shape using the backboned aTAM. In Fig. 7, we plot the difference in these upper bounds on the tile complexity for a single target shape, as a function of shape size. We consider both rectangular target shapes, and rectangles with a bulge to produce a target with lower symmetry. In this case, the Hamiltonian path taken by the backbone is assumed to zig-zag down the rectangle along each row starting from the top row of the rectangle (Figure 7 (a) and 7 (b)).
As can be seen in Figure 7, the bounds suggest that the very smallest rectangles can be constructed with fewer unique tile types using the base aTAM compared to the backboned aTAM. This fact is likely because unbackboned assembly is better able to take advantage of the symmetry present in a rectangle ([12]), whereas the backboned aTAM is an inherently asymmetric assembly method. As expected, the asymmetric bulged rectangles display no such behaviour as they have no symmetries that base aTAM can take advantage of. For larger rectangles, the backboned aTAM becomes more efficient than the base aTAM due to its ability to use the same tile types in locations with distinct symmetry, and does so well before the bound of the universal kit size is reached.
@@ 3.3 The Sequenced aTAM has No Universal Assembly Kit
We have shown that the backboned aTAM possesses a universal assembly kit, whereas the conventional temperature-1 aTAM does not. At least in part, this difference is due to the information content of the sequence. However, an additional difference is provided by the geometric constraints on growth provided by the backbone. In an attempt to disentangle the role of sequence information and geometric constraints, we consider the sequenced aTAM defined in Section 2, which is identical to the backboned aTAM, except added tiles are no longer constrained to neighbour immediately preceding tiles. We will see that no universal assembly kit exists that allows the sequenced aTAM to deterministically assemble every shape in its shape space, even when allowing for neutral faces.
Figure 8: Intuition for Lemma 3.9. (Left) A configuration $A$ given some starting configuration $A_{0}$ (white), and subconfigurations $A_{\epsilon}$ (blue) and $A_{\epsilon^{\prime}}$ (yellow). (Right) Assume the faces in purple have the same glue, then subshapes that grow from the purple faces can be flipped while maintaining the same tile type sequence $Q$, forming an alternate configuration $A^{\prime}$.
Our proof proceeds from the following intuition. For an instance of the sequenced aTAM, shapes that grow from a face of a configuration $A_{t}$ for times $t^{\prime}>t$ are encoded fully by the glue of the face and the tile type sequence past past time $t$ (if we ignore the ‘blocking’ of growth by preexisting tiles). Hence, if we have two tile faces with the same glue type at a time $t$, then (ignoring tile blocking), we cannot stop a shape that grows from one of the faces from also growing on the other face (Figure 8). Hence, if we need to grow $N$ different shapes from $N$ different faces at a time $t$, we require at least $N$ different glue types. If the growth of certain shapes requires an arbitrarily large number of open faces, $N\rightarrow\infty$, then a finite universal assembly kit will be impossible.
@@
Lemma 3.9 formalizes this intuition. Using this lemma, Theorem 3.11 in the main text builds a set of assumptions that allow us to avoid tile blocking and forces us to ‘open’ at least $N$ distinct glue types when $g\geq 0$. This result is extended to unrestricted $g$ in Theorem A.28 in the appendix. We then develop a class of shapes that obey these assumptions for arbitrarily large $N$ in Definition 3.14, and complete the proof in Theorem 3.17.
@@
Over the course of our proofs, we will be relying on sufficient and necessary conditions for some configuration $A_{t}$ to be part of a complete trajectory. We note that from the definitions of the backboned and sequenced aTAM, it is clear that $g-validity$ is a necessary condition for any configuration in a complete trajectory. Together with the consistency of starting configurations and added tiles matching the sequence $Q$ (either to the end of $Q$ or until the point of premature termination), g-validity becomes sufficient in ensuring that a given trajectory is complete. We now proceed by formalizing our intuition that open faces growing distinct subshapes require distinct faces, under some set of assumptions:
Lemma 3.9.
@@
    Let $A_{\epsilon}$ be a configuration adjacent to $A_{0}$ uniquely through face pair $(\dot{e}_{\epsilon},\ddot{e}_{\epsilon})$ ($\dot{e}_{\epsilon}$ in $A_{\epsilon}$ and $\ddot{e}_{\epsilon}$ in $A_{0}$), while $A_{\epsilon^{\prime}}$ is a configuration adjacent to $A_{0}$ uniquely through face pair $(\dot{e}_{\epsilon^{\prime}},\ddot{e}_{\epsilon^{\prime}})$ and assume $A_{\epsilon}$ and $A_{\epsilon^{\prime}}$ are not adjacent to each other. Assume further that some complete trajectory $\Psi=(A_{0},A_{1},A_{2},...,A_{\omega})$ generated by $T$ exists such that $A_{\omega}=A_{0}+A_{\epsilon}+A_{\epsilon^{\prime}}+A_{c}$ for some arbitrary configuration $A_{c}$ not adjacent to $A_{\epsilon}$ or $A_{\epsilon^{\prime}}$.
@@
    Let $A_{\epsilon\rightarrow\epsilon^{\prime}}$ be an affine transformation of $A_{\epsilon}$ that maps $\dot{e}_{\epsilon}$ to $\dot{e}_{\epsilon^{\prime}}$, and similarly for $A_{\epsilon^{\prime}\rightarrow\epsilon}$. Then, assume that $A_{\omega}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime}}+A_{\epsilon^{\prime}\rightarrow\epsilon}+A_{c}$ is $g$-valid.
Then, there exists another trajectory $\Psi^{\prime}=(A_{0}^{\prime},A_{1}^{\prime},...,A_{\omega}^{\prime},...)$ complete with respect to $\mathcal{P}$ such that the terminal configuration of Ψsuperscriptnormal-$\Psi^{\prime}$ is $A_{{\omega}}^{\prime}$ or has $A_{{\omega}}^{\prime}$ as a subconfiguration.
Proof 3.10.
We show that some selection of $A_{t}^{\prime}$ for Ψsuperscriptnormal-$\Psi^{\prime}$ meets the definition of a complete trajectory. First, by setting $A_{0}=A_{0}^{\prime}$, starting configuration consistency is established. Configurations in the trajectory $\Psi$ can be decomposed as $A_{t}=A_{0}+A_{\epsilon,t}+A_{\epsilon^{\prime},t}+A_{c,t}$, where $A_{x,t}$ is a subconfiguration of $A_{x}$ (present at time t). We require that it is possible to construct a trajectory Ψsuperscriptnormal-$\Psi^{\prime}$ complete with respect to $T$, such that each entry $A_{t}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime},t}+A_{\epsilon^{\prime}\rightarrow\epsilon,t}+A_{c,t}$ up to $t=\omega$ is g-valid. This is always true at $t=0$, and tile addition at time $t+1$ (following the tile type sequence $Q$) can always result in a g-valid $A_{t+1}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime},t+1}+A_{\epsilon^{\prime}\rightarrow\epsilon,t+1}+A_{c,t+1}$ up to $t+1=\omega$. Since the order of tiles added also matches the sequence $Q$, the condition for complete trajectories of squence aTAM is obeyed, and the theorem is thus true by induction.

### v2

@@ 2 Model Definitions
    The trajectory terminates upon reaching $A_{\zeta}$ or when no such single tile configuration $a_{t}$ can be added. In the latter case, the trajectory is said to have been prematurely terminated.
By this definition, the first tile added in an empty configuration can be placed at any coordinate, in any orientation. The set of assembled configurations for an instance of the backboned aTAM is defined as the set containing all terminal configurations for all complete trajectories generated by the backboned aTAM instance. The set of assembled shapes for a backboned aTAM instance is thus the set containing all shapes of assembled configurations of the backboned aTAM instance.
We further define a model that we call the sequenced aTAM, which retains the sequence of the backboned aTAM, but not the backbone constraints. The definition of a sequenced aTAM instance, along with the definition of its complete trajectories, are as follows. A sequenced aTAM instance is a 3-tuple $(\mathcal{A},Q,g)$ (with $\mathcal{A},Q,g$ defined identically to a backboned aTAM instance) that generates a set of complete trajectories obeying Definition 3, rather than Definition 2.
Definition 3.
Consider a sequenced aTAM instance $(\mathcal{A},Q,g)$, with the sequence of tile types $Q=(\Theta_{1},\Theta_{2},..,\Theta_{\zeta})$. A trajectory $\Psi=(A_{0},A_{1},A_{2},..)$ is said to be complete with respect to $(\mathcal{A},Q,g)$ if the following hold.
  1. 1.
    Starting configurations are consistent, that is $A_{0}=\mathcal{A}$.
  2. 2.
    $A_{t}=A_{t-1}+a_{t}$ with $a_{t}$ chosen randomly under the constraints:
    1. (a)
      $A_{t}\neq\infty$ for any $t$.
    2. (b)
    3. (c)
      Any cut of $G_{A}(A_{t})$ is such that the sum of interaction strengths over cut-edges $\sum_{i}g(\sigma_{i},\sigma_{i}^{\prime})>0$, where $\sigma_{i}$ and $\sigma_{i}^{\prime}$ are neighbouring faces corresponding to an edge. A configuration $A_{t}$ that obeys this assumption for some interaction function $g$ is called g-valid.
  3. 3.
    The trajectory terminates upon reaching $A_{\zeta}$ or when no such single tile configuration $a_{t}$ can be added. In the latter case, the trajectory is said to have been prematurely terminated.
The key difference in the definition of a complete trajectory for the sequenced aTAM relative to the backboned aTAM is that it does not mandate that new tiles are placed adjacent to the previous tile. This difference allows for the same $\mathcal{A},Q,g$ to generate a distinct set of trajectories
@@ 3 Results
To address the question of whether various aTAM models can efficiently encode the deterministic assembly of shapes into their sequences, we consider the existence of universal assembly kits, first for the backboned aTAM, and eventually for the sequenced aTAM. A universal assembly kit is defined as a finite tile type set and corresponding strength function that allows the deterministic assembly of any finite shape within the shape space of the model. Here, the starting configuration must be the empty set and the sequence of tiles can be arbitrary, within those sequences permitted by the finite tile set.
3.1 Universal Assembly Kits for Backboned aTAM
We begin our results with the following Lemma.
@@ 3.1 Universal Assembly Kits for Backboned aTAM
Proof.
This lemma trivially proceeds from definition 2, as added tiles must neighbour the last added tile. ∎
Theorem 1.
A universal assembly kit for the backboned aTAM exists.
Proof.
Since any shape that can be assembled by a backboned aTAM system is a shape of some self-avoiding walk, we can always assemble any shape that is achievable by a backboned aTAM system by encoding the set of left, right or forward moves of the underlying self-avoiding walk. We use glues $0,1$ and $2$, along with the following strength function:
$g=\left\{\begin{array}[]{ll}1,&\mbox{if}\ (\sigma,\sigma^{\prime})=(1,2)\mbox{or }(\sigma,\sigma^{\prime})=(2,1),\\ 0,&\mbox{otherwise}.\end{array}\right.$ (1)
Then, each left, right or forward move can be performed by a specific “directed” tile, shown in Figure 3. An additional tile encoding the start of the shape is required, but there is no need for an end tile as any of the directed tiles can be placed at the end of any self-avoiding walk without impacting the final shape. Hence, a universal assembly kit with 4 tiles is sufficient to assemble any shape achievable by the backboned aTAM, without requiring a starting seed configuration or needing to scale up the shape. ∎
Given this theorem, the following corollary trivially holds.
Corollary 1.
Proof.
Any self-avoiding path can be constructed using the method in Theorem 1. ∎
@@
Figure 4: A figure illustrating the difficulties associated with assembly with only attractive (i.e. without neutral) inter-tile interactions. While in principle one can replace the neutral interface type 00 with a self-attractive interface type $3$, such attractive interfaces can pull tiles towards unintended positions (right).
These initial results are fairly straightforward. However, the construct we have arrived at is somewhat artificial when considering the physical system being emulated (co-translational folding), since the shape is completely determined by the backbone routing and non-backbone-adjacent interactions are irrelevant. This result is only possible due to the use of neutral interactions, or by allowing adjacent tiles to have weak repulsive interactions (that are not sufficiently strong to prevent the addition of a tile). We therefore now consider the backboned aTAM with an added restriction that all inter-tile interactions in any configuration are attractive. There is a sense in which this variant is closer to biological folding maps, where matching non-backbone interactions are necessary for accurate folding. One strategy for devising a universal assembly kit is to use the left-right-forward tile types in 3, but to replace the 00 glues with custom glues depending on interactions. Immediately from Figure 4, more glue types, and hence tile types, would be required to avoid incorrect assembly. We now proceed to identify a set of glues and tile types that can assemble all possible shapes, thereby proving the existence of a universal assembly kit.
Theorem 2.
@@
Proof.
The existence of some infinite tile set is trivial, as in the construction of all possible shapes, one can use the tile set in Theorem 1, but replace all glue types of all faces that neighbour another tile face with some unique attractive glue type. However,a universal assembly kit requires a finite number of tiles and reusing the same attractive glue type multiples times within a configuration can result in non-deterministic assembly as tiles can be attracted into incorrect positions (Figure 4). Hence, we must proceed to construct a scheme for numbering tile types such that only a finite number of glue types $N$ are required while still guaranteeing deterministic shape assembly. Let the glue types be labeled $\{0,1,2,3,4,5...N-1\}$, and let the strength function $g$ be defined as follows.
$g(\sigma,\sigma^{\prime})=\left\{\begin{array}[]{ll}1,&\mbox{if}\ (\sigma,\sigma^{\prime})=(1,2)\mbox{ or }(\sigma,\sigma^{\prime})=(2,1)\mbox{ or }(\sigma=\sigma^{\prime}\mbox{ and }\sigma>2),\\ -3,&\mbox{otherwise}.\end{array}\right.$ (2)
@@
These tile types are constructed by analogy with the tile types used in the proof of Theorem 1. For example, $\hat{\Theta}_{0}$ is a set of tile types with glue types 1 and 2 (backbone glue types) in the same pattern as tile type 0 in Fig. 3, but with variable glues on the other faces; $\hat{\Theta}_{0}(\sigma,\sigma^{\prime})$ identifies a specific member of that set. Non-backbone glue types are divided into zero and non-zero types; we denote by $B^{c}({\Theta})$ the set of non-zero, non-backbone glue types on tile type ${\Theta}$. Tile faces endowed with backbone and non-backbone glue types are correspondingly called backbone faces and non-backbone faces, respectively.
@@
We present an algorithm for selecting a sequence $Q$, using a subset of the tiles from Definition 4 with $N=7$, to define a backboned aTAM system $(\mathcal{A},Q,g)$ that can deterministically assemble any given shape within the shape space of the backboned aTAM. The interacting directed tile types in Definition 4, along with the strength function in Equation 2, therefore define a universal assembly kit for the backboned aTAM, proving the existence of such a kit by construction.
Let $\Psi=(A_{\rm empty},A_{1},...A_{\zeta})$ be a complete trajectory that assembles a shape $\overline{S}$, defining an arbitrary Hamiltonian path through the shape. A sequence of backbone faces consistent with this trajectory can then be selected as in the proof of Theorem 1, fixing whether each tile in the sequence is drawn from the set $\hat{\Theta}_{T}$, $\hat{\Theta}_{0}$, $\hat{\Theta}_{1}$, $\hat{\Theta}_{3}$ or $\hat{\Theta}_{H}$. Hence, we only need to select the non-backbone interactions of tile types in the sequence in such a way as to ensure deterministic production of the desired shape.
@@
    Non-backbone faces of $a_{t}$ with no neighbour in $A_{\zeta}$ (the terminal configuration of $\Psi$) are assigned the repulsive glue type 0.
@@
    Non-backbone faces of $a_{t}$ with no neighbour in $A_{t}$ but with a neighbour in $A_{\zeta}$ are assigned an unknown glue type $\sigma_{u}$
@@
Figure 6: An illustration of the inequalities on the faces of tiles when constructing an assembly kit for the backboned aTAM with only attractive inter-tile interactions. After setting the glue types of known faces, there remains one face with unknown glue type $\sigma_{u}$. $\sigma_{u}\neq 3$, $\sigma_{u}\neq 5$ and $\sigma_{u}\neq 6$, so $\sigma_{u}=4$ is the only correct option if we restrict ourselves to $4$ attractive non-backbone glue types.
@@
The arguments above break down for the penultimate tile as the final tile can have three non-backbone-connected neighbours. However, in this case all of the final neighbours of $a_{\zeta}$ and $a_{\zeta-1}$ are already in place. Therefore the non-backbone glue types of $a_{\zeta-1}$ are either specified by neighbouring tiles that are present in the configuration $A_{\zeta-1}$, or 0. They cannot, therefore, cause any ambiguity when $a_{\zeta}$ is placed, with glue types set either to match neighbours present in configuration $A_{\zeta}$, or 0. Taking $N=7$, 208 tiles are defined by Definition 4. This number provides an upper bound on the size of the minimal universal assembly kit for the backboned aTAM.
3.2 Production of Specific Shapes with the Backboned aTAM
Having derived universal assembly kits for different variants of the backboned aTAM, we now consider how different assembly models perform when assembling specific shapes. In doing so, we hope to further verify our intuition that sequence information facilitates the assembly of shapes with fewer distinct tile types, and also observe differences in the performances of different sequence-directed assembly models. We consider the following three assembly models in this section.
  1. 1.
    Backboned aTAM with arbitrary interactions (i.e., allowing neutral inter-tile interactions).
  2. 2.
    Backboned aTAM with only attractive inter-tile interactions.
  3. 3.
    Sequence-free polyomino assembly (temperature-1 aTAM with tile rotation).
@@ 3.2 Production of Specific Shapes with the Backboned aTAM
Figure 7: We consider the construction of rectangles (a) and bulged rectangles (b) with $m$ rows and $n$ columns. The backbone path for the backboned aTAM is given in purple. The bulge location is fixed to the second and third rows of the rightmost column.
Figure 8: Upper bounds on the tile complexity, or the number of tiles needed to construct (a) rectangles and (b) bulged rectangles of various row and column numbers using different assembly maps. Backboned maps tend to be capped in the number of tiles they require for arbitrarily large shapes (of the given types), while the sequence-free assembly map requires an unbounded number of tile types. Forcing attractive inter-tile interactions tends to increase the number of tile types required.
@@
For these models, we consider upper bounds on the minimum number of tiles, also called the tile complexity, needed to construct two types of target shape, rectangles and rectangles with a bulge (Figures 7(a) and 7(b)), starting from the empty configuration. We apply the method in (?) to upper bound a minimal tile complexity for a given shape using sequence-free polyomino assembly. For the backboned aTAM with arbitrary interactions, an upper bound on the minimum tile set is easy to derive, as we can follow some arbitrary hamiltonian path and use the tile types in Theorem 1. For the backboned aTAM with only attractive inter-tile interactions, we can follow a face assignment algorithm implied by Theorem 2 to find an upper bound on the tile complexity. For these assembly schemes, the hamiltonian path taken is assumed to zig-zag down the rectangle along each row starting from the top row of the rectangle (Figure 8 (a) and 8 (b)).
In Fig. 8, we plot upper bounds on the minimum tile complexities of each target shape using different assembly maps. The very smallest rectangles can be constructed with fewer unique tile types using sequence-free assembly than with backbone-directed assembly, as sequence-free assembly is better able to utilize symmetries within the shape (?). As expected, the asymmetric bulged rectangles display no such behavior as they have no symmetries of which the base aTAM can take advantage. In both cases, the number of tiles required for the base aTAM increases unboundedly for large rectangles, as anticipated. By contrast, the backboned aTAM models show a growth in the number of tile types that plateaus, verifying our intuition that sequences facilitate the assembly of large shapes with a small number of building blocks.
@@ 3.3 The Sequenced aTAM has No Universal Assembly Kit
We have shown that the backboned aTAM possesses a universal assembly kit. We have also confirmed by example that, for large structures, one can achieve deterministic assembly of a single structure with smaller tile sets. At least in part, this difference arises because each instance of the backboned aTAM is associated with a sequence that acts as an information-containing program, in addition to the rules of interaction quantified by the glues. However, an additional difference is provided by the geometric constraints on growth provided by the backbone.
In an attempt to disentangle the role of sequence information and geometric constraints, we have introduced the sequenced aTAM defined in Section 2. We will now show that no universal assembly kit exists that allows the sequenced aTAM to deterministically assemble every shape in its shape space, regardless of constraints on neutral versus attractive inter-tile interactions.
Figure 9: Intuition for Lemma 2. (Left) A configuration $A$ given some starting configuration $A_{0}$ (white), and subconfigurations $A_{\epsilon}$ (blue) and $A_{\epsilon^{\prime}}$ (yellow). (Right) Assume the faces in purple have the same glue, then subshapes that grow from the purple faces can be flipped while maintaining the same tile type sequence $Q$, forming an alternate configuration $A^{\prime}$.
Our proof proceeds from the following intuition. For an instance of the sequenced aTAM, shapes that grow from a face of a configuration $A_{t}$ for times $t^{\prime}>t$ are encoded fully by the glue of the face and the tile type sequence past time $t$ (if we ignore the ‘blocking’ of growth by preexisting tiles). Hence, if we have two tile faces with the same glue type at a time $t$, then (ignoring tile blocking), we cannot stop a shape that grows from one of the faces from also growing on the other face (Figure 9). Hence, if we need to grow $N$ different shapes from $N$ different faces at a time $t$, we require at least $N$ different glue types. If the growth of certain shapes requires an arbitrarily large number of open faces, $N\rightarrow\infty$, then a finite universal assembly kit will be impossible.
@@
Lemma 2 formalizes this intuition. Using this lemma, Theorem 3 in the main text builds a set of assumptions that allow us to avoid tile blocking and forces us to ‘open’ at least $N$ distinct glue types when $g\geq 0$. This result is extended to unrestricted $g$ in Theorem 6 in the appendix. We then develop a class of shapes that obey these assumptions for arbitrarily large $N$ in Definition 6, and complete the proof in Theorem 4. Readers interested only in the consequences of this result in terms of the efficiency with which the variant aTAM models exploit information within their sequence programs may skip directly to Section 3.4.
@@
Over the course of our proofs, we will be relying on sufficient and necessary conditions for some configuration $A_{t}$ to be part of a complete trajectory. We note that, from the definitions of the backboned and sequenced aTAM, it is clear that $g-validity$ is a necessary condition for any configuration in a complete trajectory. Together with the consistency of starting configurations and added tiles matching the sequence $Q$ (either to the end of $Q$ or until the point of premature termination), g-validity becomes sufficient in ensuring that a given trajectory is complete. We now proceed by formalizing our intuition that open faces growing distinct subshapes require distinct faces, under some set of assumptions:
Lemma 2.
@@
    Let $A_{\epsilon}$ be a configuration adjacent to $A_{0}$ uniquely through face pair $(\dot{e}_{\epsilon},\ddot{e}_{\epsilon})$ ($\dot{e}_{\epsilon}$ in $A_{\epsilon}$ and $\ddot{e}_{\epsilon}$ in $A_{0}$), while $A_{\epsilon^{\prime}}$ is a configuration adjacent to $A_{0}$ uniquely through face pair $(\dot{e}_{\epsilon^{\prime}},\ddot{e}_{\epsilon^{\prime}})$ and assume $A_{\epsilon}$ and $A_{\epsilon^{\prime}}$ are not adjacent to each other. Assume further that some complete trajectory $\Psi=(A_{0},A_{1},A_{2},...,A_{\zeta})$ generated by $T$ exists such that $A_{\zeta}=A_{0}+A_{\epsilon}+A_{\epsilon^{\prime}}+A_{c}$ for some arbitrary configuration $A_{c}$ not adjacent to $A_{\epsilon}$ or $A_{\epsilon^{\prime}}$.
@@
    Let $A_{\epsilon\rightarrow\epsilon^{\prime}}$ be an affine transformation of $A_{\epsilon}$ that maps $\dot{e}_{\epsilon}$ to $\dot{e}_{\epsilon^{\prime}}$, and similarly for $A_{\epsilon^{\prime}\rightarrow\epsilon}$. Then, assume that $A_{\zeta}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime}}+A_{\epsilon^{\prime}\rightarrow\epsilon}+A_{c}$ is $g$-valid.
Then, there exists another trajectory $\Psi^{\prime}=(A_{0}^{\prime},A_{1}^{\prime},...,A_{\zeta}^{\prime},...)$ complete with respect to $\mathcal{P}$ such that the terminal configuration of $\Psi^{\prime}$ is $A_{{\zeta}}^{\prime}$ or has $A_{{\zeta}}^{\prime}$ as a subconfiguration.
Proof.
We show that some selection of $A_{t}^{\prime}$ for $\Psi^{\prime}$ meets the definition of a complete trajectory. First, by setting $A_{0}=A_{0}^{\prime}$, starting configuration consistency is established. Configurations in the trajectory $\Psi$ can be decomposed as $A_{t}=A_{0}+A_{\epsilon,t}+A_{\epsilon^{\prime},t}+A_{c,t}$, where $A_{x,t}$ is a subconfiguration of $A_{x}$ (present at time t). We require that it is possible to construct a trajectory $\Psi^{\prime}$ complete with respect to $T$, such that each entry $A_{t}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime},t}+A_{\epsilon^{\prime}\rightarrow\epsilon,t}+A_{c,t}$ up to $t=\zeta$ is g-valid. This condition is always true at $t=0$, and tile addition at time $t+1$ (following the tile type sequence $Q$) can always result in a g-valid $A_{t+1}^{\prime}=A_{0}+A_{\epsilon\rightarrow\epsilon^{\prime},t+1}+A_{\epsilon^{\prime}\rightarrow\epsilon,t+1}+A_{c,t+1}$ up to $t+1=\zeta$. Since the order of tiles added also matches the sequence $Q$, the condition for complete trajectories of sequenced aTAM is obeyed, and the theorem is thus true by induction. ∎